# Adjoint motives of modular forms and the Tamagawa number conjecture

### v1

@@ 1.2. Premotivic structures
  • is $L$-admissible everywhere if it is $L$-admissible at $p$ for all primes pp.
@@
We say that $\mathcal{M}$} has good reduction at $p$, is $L$-admissible at $p$ or is $L$-admissible everywhere according to whether the same is true for $\mathbb{Q}$$\otimes{\mathcal{M}}$. Note that if $\mathcal{M}$} is $L$-admissible at $p$ and $p$ is not invertible in $\mathcal{O}$}, then $\mathcal{M}$} necessarily has good reduction at pp.
@@
Note that if $\mathcal{M}$} is an object of ${\mathcal{P}M}_{K}^{S}$, then $\operatorname{{\mathop{\rm End}\nolimits}}{\mathcal{M}}$ is a finitely generated ${\mathcal{O}}_{K}$-module. If II is an ${\mathcal{O}}_{K}$-submodule of $\operatorname{{\mathop{\rm End}\nolimits}}{\mathcal{M}}$, then we define an object ${\mathcal{M}}[I]$ of ${\mathcal{P}M}_{K}^{S}$ as the kernel of
@@
where $x_{1},\ldots,x_{r}$ generate II. This is independent of the choice of generators. This applies in particular when II is the image in $\operatorname{{\mathop{\rm End}\nolimits}}{\mathcal{M}}$ of an ideal in a commutative ${\mathcal{O}}_{K}$-algebra $R$ mapping to $\operatorname{{\mathop{\rm End}\nolimits}}{\mathcal{M}}$, or the augmentation ideal in ${\mathcal{O}}_{K}[G]$ where $G$ is a group acting on $\mathcal{M}$}. In the latter case, we write ${\mathcal{M}}^{G}$ instead of ${\mathcal{M}}[I]$.
@@ 1.3. Basic examples
Suppose that we are given a character $\psi:\hat{{\mathbb{Z}}}^{\times}\to K^{\times}$. We regard $\psi$ also as a character of $\mathbb{A}$$^{\times}$ and $G_{\mathbb{Q}}$ via the isomorphisms $\hat{{\mathbb{Z}}}^{\times}\cong{\mathbb{A}}^{\times}/{\mathbb{R}}_{>0}^{\times}{\mathbb{Q}}^{\times}\cong G_{\mathbb{Q}}^{\mathop{\rm ab}\nolimits}$, where the first isomorphism is induced by the natural inclusion $\hat{{\mathbb{Z}}}^{\times}\to{\mathbb{A}}^{\times}$ and the second is given by class field theory. (Our convention is that a uniformizer in ${\mathbb{Q}}_{p}^{\times}$ maps to $\operatorname{{\mathop{\rm Frob}\nolimits}}_{p}$ in the Galois group of any abelian extension of $\mathbb{Q}$} unramified at pp.)
Suppose we are given an embedding of a number field $F\to\bar{{\mathbb{Q}}}$ such that $\psi$ is trivial on the image of $G_{F}$ in $G_{\mathbb{Q}}$. We can then regard $\psi$ as a character of $G=\operatorname{{\mathop{\rm Gal}\nolimits}}(F/{\mathbb{Q}})$ and define the Dirichlet premotivic structure $M_{\psi}$ as $(V\otimes M_{F})^{G}$ where $V=K$ with $G$ acting by $\psi$. One can check that the construction is independent of the choice of $F$ and embedding $F\to\bar{{\mathbb{Q}}}$ as above. To describe $M_{\psi}$ explicitly, we choose $F={\mathbb{Q}}(e^{2\pi i/N})\subset{\mathbb{C}}$ where $\psi$ has conductor NN. We let $\tau_{0}:F\to\bar{{\mathbb{Q}}}$ denote the embedding compatible with our fixed $\bar{{\mathbb{Q}}}\to{\mathbb{C}}$, and we regard $\psi$ as a Dirichlet character via the canonical isomorphism $({\mathbb{Z}}/N{\mathbb{Z}})^{\times}\cong G$. Let $S$ denote the set of primes in $K$ lying over those dividing NN and define an object ${\mathcal{M}}_{\psi}$ of ${\mathcal{P}M}_{K}^{S}$ by $({\mathcal{V}}\otimes{\mathcal{M}}_{F})^{G}$ where ${\mathcal{V}}={\mathcal{O}}_{K}^{S}$ with $G$ acting by $\psi$. We then have:
@@ 2. Premotivic structures for modular forms
In this section we review the construction of premotivic structures associated to the space of modular forms of weight $k$ and level NN. More precisely, if $k\geq 2$ and $N\geq 3$, we construct objects of ${\mathcal{P}M}_{\mathbb{Q}}^{S}$ whose de Rham realization contains the space of such forms, where $S=S_{N}$ is the set of primes dividing $Nk!$.
2.1. Level NN modular curves
Let $k$ and NN be integers with $k\geq 2$ and $N\geq 3$. Let $T=\operatorname{{\mathop{\rm Spec}\nolimits}}{\mathbb{Z}}[1/Nk!]$, and consider the functor which associates to a $T$-scheme $T^{\prime}$ the set of isomorphism classes of generalized elliptic curves over $T^{\prime}$ with level NN structure [D-R, IV.6.6]. By [D-R, IV.6.7], the functor is represented by a smooth, proper curve over $T$. We denote this curve by $\bar{X}$, and we let $\bar{s}:\bar{E}\to\bar{X}$ denote the universal generalized elliptic curve with level NN structure. We let $X$ denote the open subscheme of $\bar{X}$ over which $\bar{E}$ is smooth. Then $X$ is the complement of a reduced divisor, called the cuspidal divisor, which we denote by $X^{\infty}$. We let $E=\bar{s}^{-1}X$, $s=\bar{s}|_{E}$ and $E^{\infty}=\bar{E}\times_{\bar{X}}X^{\infty}$. Using the arguments of [D-R, VII.2.4], one can check that $\bar{E}$ is smooth over $T$ and $E^{\infty}$ is a reduced divisor with strict normal crossings (in the sense of [G-M, 1.8] as well as [SGA1, XIII.2.1]).
@@ 2.1. Level NN modular curves
where $(m,n)\in{\mathbb{Z}}\times{\mathbb{Z}}$ acts on $\mathfrak{H}$$\times{\mathbb{C}}$ via $(\tau,z)\mapsto(\tau,z+m\tau+n)$, and $\gamma={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$}}\in\Gamma(N)$ acts by sending the class of $(\tau,z)$ to that of $(\gamma(\tau),(c\tau+d)^{-1}z)$. We can regard $E_{N}=\coprod E_{N,t}$ as a complex analytic family of elliptic curves over $X_{N}$. We can then extend $E_{N}$ to a family $\bar{E}_{N}$ of generalized elliptic curves over $\bar{X}_{N}$ using analytic Tate curves, as in [D-R, VII.4]. More precisely, for $\alpha>0$, we let $D_{\alpha}$ denote the disk in $D[q^{1/N}]$ defined by $|q|<e^{-2\pi\alpha}$ and let $D_{\alpha}^{0}$ denote the corresponding punctured disk. Then over a punctured neighborhood around each cusp in $\bar{X}_{N}$, we have an isomorphism between $E_{N}$ and the NN-sided analytic Tate curve $E_{q}^{\mathop{\rm an}\nolimits}$ over $D_{1}^{0}$, and we define $\bar{E}_{N}$ by gluing NN-sided analytic Tate curves over $D_{1}$ to $E_{N}$. For each $a$, we define a level NN-structure on $E_{N,t}$ by the pair of sections $X_{N,t}\to E_{N,t}$ induced by $\tau\mapsto(\tau,\tau/N)$ and $\tau\mapsto(\tau,t/N)$. The resulting level NN-structure on $E_{N}$ in fact extends to $\bar{E}_{N}$.
Using for example [Ha, App. B, 3.4], one checks that $\bar{E}_{N}$ is algebraic, so $\bar{E}_{N}\to\bar{X}_{N}$ is the analytification of a generalized elliptic curve $\bar{E}_{N}^{\mathop{\rm alg}\nolimits}\to\bar{X}_{N}^{\mathop{\rm alg}\nolimits}$. One can check the resulting morphism $\bar{X}_{N}^{\mathop{\rm alg}\nolimits}\to\bar{X}_{\mathbb{C}}$ induces a bijection $X_{N}\to X({\mathbb{C}})$, and since both are smooth, proper algebraic curves over $\mathbb{C}$}, the map is an isomorphism. The analytification of the universal generalized elliptic curve with level NN-structure is therefore isomorphic to $\bar{E}_{N}\to\bar{X}_{N}$ with the level NN-structure defined above.
$2.$ Realizations of level NN modular forms
@@
Over a neighborhood of each cusp in $\bar{X}^{\mathop{\rm an}\nolimits}$, $\bar{E}^{\mathop{\rm an}\nolimits}$ is isomorphic to the NN-sided analytic Tate curve $E_{q}^{\mathop{\rm an}\nolimits}$ over $D_{1}$. We claim that there is an isomorphism between the pull-back of ${\mathcal{F}}_{{\mathop{\rm dR}\nolimits}}^{k,{\mathop{\rm an}\nolimits}}$ to $D_{1}$ and ${\mathcal{O}}_{D_{1}}^{k-1}$ such that the corresponding connection on ${\mathcal{O}}_{D_{1}}^{k-1}$ is given by $\nabla(f)=(A(t)f+tf^{\prime})\dfrac{dt}{t}$ where $A(t)\in M_{k-1}({\mathcal{O}}_{D_{1}})$ is such that $A(0)$ is nilpotent. Indeed this follows from the case $k=3$ which follows from the calculation in Appendix 1.3 of [Katz2] for the 1-sided Tate curve.
@@
In the case of $Y=\bar{Y}=\operatorname{{\mathop{\rm Spec}\nolimits}}{\mathcal{O}}_{F}$ where $F$ is an unramified extension of $\mathbb{Q}$$_{\ell}$, Faltings’ category can be identified with the full subcategory of ${{\mathcal{M}\mathcal{F}}}^{0}_{\mathop{\rm tor}\nolimits}({\mathcal{O}}_{F})$ whose objects AA satisfy $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a+1}A=0$. Furthermore, identifying the category of étale sheaves on $\operatorname{{\mathop{\rm Spec}\nolimits}}F$ with that of continuous $G_{F}$-modules, one finds that the functor $\mathbb{D}$} agrees with $\underline{U}_{S}$ (see [Fo3, 3.8]).
@@
2.3. Level NN modular forms
@@ 2.5. Weight filtrations and summary
where $V=\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}({\mathbb{C}}\otimes M_{c,{\mathop{\rm dR}\nolimits}})\cong\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}W_{k-1}({\mathbb{C}}\otimes M_{\mathop{\rm dR}\nolimits})$, and this is immediate from Lemma $2.$
@@
We now summarize main properties of the premotivic structures ${\mathcal{M}}_{N}$ and $M_{N}$ for level NN modular forms.
@@ 3.2. Action on curves
Suppose now that $(h,N,N^{\prime})$ is an admissible triple. Let $\bar{E}/\bar{X}$ (respectively, $\bar{E}^{\prime}/\bar{X}^{\prime}$) denote the universal generalized elliptic curve with level NN (respectively $N^{\prime}$) structure.
@@
and so gives rise to a level NN-structure on $E^{\prime}/G$.
Letting $(\bar{E}^{\prime}/G)_{\rm cont}$ denote the contraction of $\bar{E}^{\prime}/G$ whose cuspidal fibers are NN-gons [D-R, IV.1], we obtain a level NN structure on $(\bar{E}^{\prime}/G)_{\rm cont}\to\bar{X}^{\prime}$. By the universal property of $\bar{E}_{/\bar{X}}$, this defines a map $\chi_{h}=\chi_{h,N,N^{\prime}}:\bar{X}^{\prime}\to\bar{X}$ such that there is an isomorphism $(\bar{E}^{\prime}/G)_{\rm cont}\to\bar{E}\times_{\bar{X}}\bar{X}^{\prime}$ of generalized elliptic curves with level NN-structures. We let $\varepsilon_{h}=\varepsilon_{h,N,N^{\prime}}$ denote the composite of the natural map $\bar{E}^{\prime}\to(\bar{E}^{\prime}/G)_{\rm cont}$ with this isomorphism, so we obtain a commutative diagram
@@
induces an isomorphism $E_{\tau^{\prime}}/G_{\tau^{\prime}}\cong E_{\tau}$ compatible with level NN-structures.
@@
Furthermore, since multiplication by $u$ is the identity on $({\mathbb{Z}}/N{\mathbb{Z}})^{2}$, the level NN-structure on $E_{\tau^{\prime}}/G_{\tau^{\prime}}$ is given by
@@
$3.$ Action on the premotivic structures
Suppose that $(h,N,N^{\prime})$ is an admissible triple. Let $S$ denote the set of primes dividing $Nk!$ and $S^{\prime}$ the set of primes dividing $N^{\prime}k!$. We let $\mathcal{M}$}, ${\mathcal{M}}_{c}$ and $\mathcal{M}$$_{!}$ (respectively, $\mathcal{M}$$^{\prime}$, ${\mathcal{M}}^{\prime}_{c}$ and $\mathcal{M}$$^{\prime}_{!}$) denote the objects of ${\mathcal{P}M}_{\mathbb{Q}}^{S}$ (respectively, ${\mathcal{P}M}_{\mathbb{Q}}^{S^{\prime}}$) structures defined in §2.5 associated to modular forms of weight $k$ and level NN (respectively, level $N^{\prime}$). We shall define compatible morphisms ${\mathcal{M}}_{c}^{S^{\prime}}\to{\mathcal{M}}_{c}^{\prime}$ and ${\mathcal{M}}^{S^{\prime}}\to{\mathcal{M}}^{\prime}$, giving also ${\mathcal{M}}_{!}^{S^{\prime}}\to{\mathcal{M}}_{!}^{\prime}$.

### v2

@@ 1.2. Premotivic structures
  • is $L$-admissible everywhere if it is $L$-admissible at $p$ for all primes $p$.
@@
We say that $\mathcal{M}$} has good reduction at $p$, is $L$-admissible at $p$ or is $L$-admissible everywhere according to whether the same is true for $\mathbb{Q}$$\otimes{\mathcal{M}}$. Note that if $\mathcal{M}$} is $L$-admissible at $p$ and $p$ is not invertible in $\mathcal{O}$}, then $\mathcal{M}$} necessarily has good reduction at $p$.
@@
Note that if $\mathcal{M}$} is an object of ${\mathcal{P}M}_{K}^{S}$, then $\operatorname{{\mathop{\rm End}\nolimits}}{\mathcal{M}}$ is a finitely generated ${\mathcal{O}}_{K}$-module. If $I$ is an ${\mathcal{O}}_{K}$-submodule of $\operatorname{{\mathop{\rm End}\nolimits}}{\mathcal{M}}$, then we define an object ${\mathcal{M}}[I]$ of ${\mathcal{P}M}_{K}^{S}$ as the kernel of
@@
where $x_{1},\ldots,x_{r}$ generate $I$. This is independent of the choice of generators. This applies in particular when $I$ is the image in $\operatorname{{\mathop{\rm End}\nolimits}}{\mathcal{M}}$ of an ideal in a commutative ${\mathcal{O}}_{K}$-algebra $R$ mapping to $\operatorname{{\mathop{\rm End}\nolimits}}{\mathcal{M}}$, or the augmentation ideal in ${\mathcal{O}}_{K}[G]$ where $G$ is a group acting on $\mathcal{M}$}. In the latter case, we write ${\mathcal{M}}^{G}$ instead of ${\mathcal{M}}[I]$.
@@ 1.3. Basic examples
Suppose that we are given a character $\psi:\hat{{\mathbb{Z}}}^{\times}\to K^{\times}$. We regard $\psi$ also as a character of $\mathbb{A}$$^{\times}$ and $G_{\mathbb{Q}}$ via the isomorphisms $\hat{{\mathbb{Z}}}^{\times}\cong{\mathbb{A}}^{\times}/{\mathbb{R}}_{>0}^{\times}{\mathbb{Q}}^{\times}\cong G_{\mathbb{Q}}^{\mathop{\rm ab}\nolimits}$, where the first isomorphism is induced by the natural inclusion $\hat{{\mathbb{Z}}}^{\times}\to{\mathbb{A}}^{\times}$ and the second is given by class field theory. (Our convention is that a uniformizer in ${\mathbb{Q}}_{p}^{\times}$ maps to $\operatorname{{\mathop{\rm Frob}\nolimits}}_{p}$ in the Galois group of any abelian extension of $\mathbb{Q}$} unramified at $p$.)
Suppose we are given an embedding of a number field $F\to\bar{{\mathbb{Q}}}$ such that $\psi$ is trivial on the image of $G_{F}$ in $G_{\mathbb{Q}}$. We can then regard $\psi$ as a character of $G=\operatorname{{\mathop{\rm Gal}\nolimits}}(F/{\mathbb{Q}})$ and define the Dirichlet premotivic structure $M_{\psi}$ as $(V\otimes M_{F})^{G}$ where $V=K$ with $G$ acting by $\psi$. One can check that the construction is independent of the choice of $F$ and embedding $F\to\bar{{\mathbb{Q}}}$ as above. To describe $M_{\psi}$ explicitly, we choose $F={\mathbb{Q}}(e^{2\pi i/N})\subset{\mathbb{C}}$ where $\psi$ has conductor $N$. We let $\tau_{0}:F\to\bar{{\mathbb{Q}}}$ denote the embedding compatible with our fixed $\bar{{\mathbb{Q}}}\to{\mathbb{C}}$, and we regard $\psi$ as a Dirichlet character via the canonical isomorphism $({\mathbb{Z}}/N{\mathbb{Z}})^{\times}\cong G$. Let $S$ denote the set of primes in $K$ lying over those dividing $N$ and define an object ${\mathcal{M}}_{\psi}$ of ${\mathcal{P}M}_{K}^{S}$ by $({\mathcal{V}}\otimes{\mathcal{M}}_{F})^{G}$ where ${\mathcal{V}}={\mathcal{O}}_{K}^{S}$ with $G$ acting by $\psi$. We then have:
@@ 2. Premotivic structures for modular forms
In this section we review the construction of premotivic structures associated to the space of modular forms of weight $k$ and level $N$. More precisely, if $k\geq 2$ and $N\geq 3$, we construct objects of ${\mathcal{P}M}_{\mathbb{Q}}^{S}$ whose de Rham realization contains the space of such forms, where $S=S_{N}$ is the set of primes dividing $Nk!$.
2.1. Level $N$ modular curves
Let $k$ and $N$ be integers with $k\geq 2$ and $N\geq 3$. Let $T=\operatorname{{\mathop{\rm Spec}\nolimits}}{\mathbb{Z}}[1/Nk!]$, and consider the functor which associates to a $T$-scheme $T^{\prime}$ the set of isomorphism classes of generalized elliptic curves over $T^{\prime}$ with level $N$ structure [D-R, IV.6.6]. By [D-R, IV.6.7], the functor is represented by a smooth, proper curve over $T$. We denote this curve by $\bar{X}$, and we let $\bar{s}:\bar{E}\to\bar{X}$ denote the universal generalized elliptic curve with level $N$ structure. We let $X$ denote the open subscheme of $\bar{X}$ over which $\bar{E}$ is smooth. Then $X$ is the complement of a reduced divisor, called the cuspidal divisor, which we denote by $X^{\infty}$. We let $E=\bar{s}^{-1}X$, $s=\bar{s}|_{E}$ and $E^{\infty}=\bar{E}\times_{\bar{X}}X^{\infty}$. Using the arguments of [D-R, VII.2.4], one can check that $\bar{E}$ is smooth over $T$ and $E^{\infty}$ is a reduced divisor with strict normal crossings (in the sense of [G-M, 1.8] as well as [SGA1, XIII.2.1]).
@@ 2.1. Level $N$ modular curves
where $(m,n)\in{\mathbb{Z}}\times{\mathbb{Z}}$ acts on $\mathfrak{H}$$\times{\mathbb{C}}$ via $(\tau,z)\mapsto(\tau,z+m\tau+n)$, and $\gamma={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$}}\in\Gamma(N)$ acts by sending the class of $(\tau,z)$ to that of $(\gamma(\tau),(c\tau+d)^{-1}z)$. We can regard $E_{N}=\coprod E_{N,t}$ as a complex analytic family of elliptic curves over $X_{N}$. We can then extend $E_{N}$ to a family $\bar{E}_{N}$ of generalized elliptic curves over $\bar{X}_{N}$ using analytic Tate curves, as in [D-R, VII.4]. More precisely, for $\alpha>0$, we let $D_{\alpha}$ denote the disk in $D[q^{1/N}]$ defined by $|q|<e^{-2\pi\alpha}$ and let $D_{\alpha}^{0}$ denote the corresponding punctured disk. Then over a punctured neighborhood around each cusp in $\bar{X}_{N}$, we have an isomorphism between $E_{N}$ and the $N$-sided analytic Tate curve $E_{q}^{\mathop{\rm an}\nolimits}$ over $D_{1}^{0}$, and we define $\bar{E}_{N}$ by gluing $N$-sided analytic Tate curves over $D_{1}$ to $E_{N}$. For each $a$, we define a level $N$-structure on $E_{N,t}$ by the pair of sections $X_{N,t}\to E_{N,t}$ induced by $\tau\mapsto(\tau,\tau/N)$ and $\tau\mapsto(\tau,t/N)$. The resulting level $N$-structure on $E_{N}$ in fact extends to $\bar{E}_{N}$.
Using for example [Ha, App. B, 3.4], one checks that $\bar{E}_{N}$ is algebraic, so $\bar{E}_{N}\to\bar{X}_{N}$ is the analytification of a generalized elliptic curve $\bar{E}_{N}^{\mathop{\rm alg}\nolimits}\to\bar{X}_{N}^{\mathop{\rm alg}\nolimits}$. One can check the resulting morphism $\bar{X}_{N}^{\mathop{\rm alg}\nolimits}\to\bar{X}_{\mathbb{C}}$ induces a bijection $X_{N}\to X({\mathbb{C}})$, and since both are smooth, proper algebraic curves over $\mathbb{C}$}, the map is an isomorphism. The analytification of the universal generalized elliptic curve with level $N$-structure is therefore isomorphic to $\bar{E}_{N}\to\bar{X}_{N}$ with the level $N$-structure defined above.
2.2. Realizations of level $N$ modular forms
@@
Over a neighborhood of each cusp in $\bar{X}^{\mathop{\rm an}\nolimits}$, $\bar{E}^{\mathop{\rm an}\nolimits}$ is isomorphic to the $N$-sided analytic Tate curve $E_{q}^{\mathop{\rm an}\nolimits}$ over $D_{1}$. We claim that there is an isomorphism between the pull-back of ${\mathcal{F}}_{{\mathop{\rm dR}\nolimits}}^{k,{\mathop{\rm an}\nolimits}}$ to $D_{1}$ and ${\mathcal{O}}_{D_{1}}^{k-1}$ such that the corresponding connection on ${\mathcal{O}}_{D_{1}}^{k-1}$ is given by $\nabla(f)=(A(t)f+tf^{\prime})\dfrac{dt}{t}$ where $A(t)\in M_{k-1}({\mathcal{O}}_{D_{1}})$ is such that $A(0)$ is nilpotent. Indeed this follows from the case $k=3$ which follows from the calculation in Appendix 1.3 of [Katz2] for the 1-sided Tate curve.
@@
In the case of $Y=\bar{Y}=\operatorname{{\mathop{\rm Spec}\nolimits}}{\mathcal{O}}_{F}$ where $F$ is an unramified extension of $\mathbb{Q}$$_{\ell}$, Faltings’ category can be identified with the full subcategory of ${{\mathcal{M}\mathcal{F}}}^{0}_{\mathop{\rm tor}\nolimits}({\mathcal{O}}_{F})$ whose objects $A$ satisfy $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a+1}A=0$. Furthermore, identifying the category of étale sheaves on $\operatorname{{\mathop{\rm Spec}\nolimits}}F$ with that of continuous $G_{F}$-modules, one finds that the functor $\mathbb{D}$} agrees with $\underline{U}_{S}$ (see [Fo3, 3.8]).
@@
2.3. Level $N$ modular forms
@@ 2.5. Weight filtrations and summary
where $V=\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}({\mathbb{C}}\otimes M_{c,{\mathop{\rm dR}\nolimits}})\cong\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}W_{k-1}({\mathbb{C}}\otimes M_{\mathop{\rm dR}\nolimits})$, and this is immediate from Lemma 2.2.
@@
We now summarize main properties of the premotivic structures ${\mathcal{M}}_{N}$ and $M_{N}$ for level $N$ modular forms.
@@ 3.2. Action on curves
Suppose now that $(h,N,N^{\prime})$ is an admissible triple. Let $\bar{E}/\bar{X}$ (respectively, $\bar{E}^{\prime}/\bar{X}^{\prime}$) denote the universal generalized elliptic curve with level $N$ (respectively $N^{\prime}$) structure.
@@
and so gives rise to a level $N$-structure on $E^{\prime}/G$.
Letting $(\bar{E}^{\prime}/G)_{\rm cont}$ denote the contraction of $\bar{E}^{\prime}/G$ whose cuspidal fibers are $N$-gons [D-R, IV.1], we obtain a level $N$ structure on $(\bar{E}^{\prime}/G)_{\rm cont}\to\bar{X}^{\prime}$. By the universal property of $\bar{E}_{/\bar{X}}$, this defines a map $\chi_{h}=\chi_{h,N,N^{\prime}}:\bar{X}^{\prime}\to\bar{X}$ such that there is an isomorphism $(\bar{E}^{\prime}/G)_{\rm cont}\to\bar{E}\times_{\bar{X}}\bar{X}^{\prime}$ of generalized elliptic curves with level $N$-structures. We let $\varepsilon_{h}=\varepsilon_{h,N,N^{\prime}}$ denote the composite of the natural map $\bar{E}^{\prime}\to(\bar{E}^{\prime}/G)_{\rm cont}$ with this isomorphism, so we obtain a commutative diagram
@@
induces an isomorphism $E_{\tau^{\prime}}/G_{\tau^{\prime}}\cong E_{\tau}$ compatible with level $N$-structures.
@@
Furthermore, since multiplication by $u$ is the identity on $({\mathbb{Z}}/N{\mathbb{Z}})^{2}$, the level $N$-structure on $E_{\tau^{\prime}}/G_{\tau^{\prime}}$ is given by
@@
3.3. Action on the premotivic structures
Suppose that $(h,N,N^{\prime})$ is an admissible triple. Let $S$ denote the set of primes dividing $Nk!$ and $S^{\prime}$ the set of primes dividing $N^{\prime}k!$. We let $\mathcal{M}$}, ${\mathcal{M}}_{c}$ and $\mathcal{M}$$_{!}$ (respectively, $\mathcal{M}$$^{\prime}$, ${\mathcal{M}}^{\prime}_{c}$ and $\mathcal{M}$$^{\prime}_{!}$) denote the objects of ${\mathcal{P}M}_{\mathbb{Q}}^{S}$ (respectively, ${\mathcal{P}M}_{\mathbb{Q}}^{S^{\prime}}$) structures defined in §2.5 associated to modular forms of weight $k$ and level $N$ (respectively, level $N^{\prime}$). We shall define compatible morphisms ${\mathcal{M}}_{c}^{S^{\prime}}\to{\mathcal{M}}_{c}^{\prime}$ and ${\mathcal{M}}^{S^{\prime}}\to{\mathcal{M}}^{\prime}$, giving also ${\mathcal{M}}_{!}^{S^{\prime}}\to{\mathcal{M}}_{!}^{\prime}$.